Verify that the given functions form a fundamental set of solutions of the differential equation on the indicated interval. Form the general solution x^3 (y''') + 6x^2 (y'') + 4xy' - 4y =0; x, x^-2, x^-2 (lnx), (0,infinity)?


1
Expert's answer
2021-06-07T18:38:14-0400

The functions "f_1(x)=x,f_2(x)=x^{-2},f_3(x)=x^{-2}lnx" are linearly independent if the Wronskian:

"W=\begin{vmatrix}
 f_1 & f_2&f_3 \\
 f'_1 & f'_2&f'_3\\
f''_1 & f''_2&f''_3\\
\end{vmatrix}\ne 0"


"W=\begin{vmatrix}
 x & x^{-2}&x^{-2}lnx \\
 1 & -2x^{-3}&-2x^{-3}lnx+x^{-3}\\
0 & 6x^{-4}&6x^{-4}lnx-2x^{-4}-3x^{-4}\\
\end{vmatrix}="


"=x\begin{vmatrix}
 -2x^{-3} & -2x^{-3}lnx+x^{-3} \\
 6x^{-4} & 6x^{-4}lnx-5x^{-4}
\end{vmatrix}-\begin{vmatrix}
 x^{-2} & x^{-2}lnx \\
 6x^{-4} & 6x^{-4}lnx-5x^{-4}
\end{vmatrix}="


"=x(-12x^{-7}lnx+10x^{-7}+12x^{-7}lnx-6x^{-7})-"

"-(6x^{-6}lnx-5x^{-6}-6x^{-6}lnx)=x(4x^{-7})+5x^{-6}=x^{-6}\ne 0"


So, the given functions form a fundamental set of solutions of the differential equation on the indicated interval.


The general solution:

"y(x)=c_1x+c_2x^{-2}+c_3x^{-2}lnx"